OpenStudy (anonymous):

Use addition or substitution to find the value of y for this set of equations. y = x - 2 7y - 4x = 10

OpenStudy (anonymous):

Substitution: y = x - 2 7y - 4x = 10 ------------------------------------------- 7y - 4x = 10 7(x - 2) - 4x = 10 7x - 14 - 4x = 10 . . . . .